In an attempt to deliver the parcel on time, the dispatch rider had to ride 10km 15 degrees SE, he then rode 11km 30 degrees NE and then takes a shortcut at 22km W. How many km did he ride NE and W combined?


1
Expert's answer
2022-03-13T18:47:55-0400

"d=d_1+d_2=11+22=33~km."